How do I solve?
When you blow across the top of a soda bottle, it acts like a closed pipe with a fundamental frequency of 495 Hz. If you pour 0.030 m of water into the bottle, shortening the air cavity, what is the new fundamental frequency? (Hint: Find the original length.) Speed of sound 343 m/s) (Unit Hz) ightaoo-ohs Corporation All Rights Resr
Answer #1

BottleL=v/(4*f) will be like a orgon pipe closed from one end. For that we know frequency(f),velocity(v) & Length(L) relation is

4L

L=v/(4*f)

L=342/(4*495)

After pouring water effective length,l=L-0.030

f= 342/4*(L-0.030)

f=599 Hz